# HumanVBench: Exploring Human-Centric Video Understanding Capabilities of MLLMs with Synthetic Benchmark Data

### v1

@@ 3.3.2 Synthesizing Questions and Preliminary Answers
In this part, the pipeline will synthesize questions and initial answers based on the video. Specifically, the video is processed using the bounding box track information of the target person. For the Human Emotion Recognition and Emotion Temporal Analysis tasks, which focus on facial expressions, we reconstruct the video with cropped face regions. For other tasks, we add a red bounding box to highlight the person in the original video, creating “marked videos”. Then, for specific tasks, we design prompts for targeted captioning with Video-MLLM. We find that video-MLLMs don’t always focus correctly on the person highlighted by the bounding box. Therefore, to further enforce the model’s attention on the target person, we add a visual cue for the target person, which is the person’s appearance in Label-4 of Figure 2.
@@
After acquiring task-specific captions, we can further analyze the distribution of caption attributes (such as the sentiment distribution in emotion recognition task) and filter the materials for generating questions based on the desired attribute distribution. Based on these captions, we utilize SOTA LLMs to formulate questions and preliminary answers based on the task definitions and the provided captions.
@@ 3.3.3 Answer Optimization and Distractors Generation
To obtain the most suitable answer to the question, we use three Video-MLLMs models to optimize the answer while generating three distractors. Specifically, for each video-MLLM, the model first answers based on the video and the question, resulting in candidate answers. Then use the same model to select a better answer between the candidate answers and the original optimal answer based on the question and video, updating the optimal answer. The discarded answers serve as raw material for generating distractors. This process is repeated three times, yielding one optimal answer as the correct option and three discarded answers. The three discarded answers need further differentiation from the correct answer, which is handled by an LLM. We ask the LLM to selectively modify the original discarded answers, adding task-specific disturbances to transform them into distractors for the question, ensuring that the generated options are clearly distinct from the correct answer.
By using the answers discarded by MLLMs to synthesize options, the options can be made more confusing for the large model, thereby further testing the model’s discriminative ability.
@@ 3.3.4 Manual Verification and Correction
Given the limitations of the annotation model and the models used for question-answer generation, errors in the multiple-choice questions may occur. Therefore, we perform manual validation at the end. Specifically, we shuffle the options of the generated multiple-choice question and ask verifiers to select the best answer. If the verifiers’ answers match the one automatically generated by the pipeline, we confirm that the question aligns with human cognition. If they do not match, we re-check and decide whether to change the correct answer option. If no best answer is found, we ask the verifier to write down the correct answer, which becomes the true correct answer for the question. Although the dataset generation process still requires some manual work, it is much less challenging and costly compared to starting from scratch, as the questions and answers have already been generated.
The Distractor-Included QA Synthesis Pipeline automates the tedious process of creating questions and options, transforming the role of humans from question creators to quality inspectors. This ensures the quality of the questions while significantly reducing labor costs. Detailed construction methods for all questions are provided in the appendix.
Moreover, in both the Human-Centric Video Annotation Pipeline and the Distractor-Included QA Generation Pipeline, all the models we use are based on the most advanced open-source models available. As more powerful and specialized models emerge, updating our pipeline with these models can further improve the quality of annotations and multiple-choice question generation.
@@ 4.1 Experimental Settings
We meticulously select 15 SOTA open-source video MLLMs. These included both visual-only models such as Chat-UniVi [19], CogVLM2-video [15], LLaVA-One-Vision [22], PLLaVA-7b [47], ShareGPT4Video [7], Otter-V [21], VILA [29], and VideoChat [23], and audio-visual models capable of analyzing both visual and audio inputs, such as the Video-LLaMA series [51, 10] and generalist MLLMs like ImageBind-LLM [13], ChatBridge [54], and OneLLM [14]. We also evaluated commercial models GPT-4o [35] and Gemini-1.5-Pro [41]. All QAs were framed as multiple-choice questions (N choose 1, with N varying across different test samples), reporting both accuracy and the performance of random guesses and graduate-level humans for reference. More implementation details can be found in the Appendix.
@@ 4.2 Main Results
Table 1 summarizes the performance evaluation results for our benchmark, revealing several key insights:
Performance of Open-Source Video-MLLMs:
There is still significant room for improvement, especially in speech-visual alignment tasks that humans excel at, which current open-source audio-visual MLLMs handle at nearly random levels. Notably, LLaVA-One-Vision emerged as the leader in visual recognition tasks but achieved a modest mean accuracy of 50.0% across 13 tasks, significantly lower than human comprehension benchmarks (86.1%). For Cross-Modal Speech-Visual Alignment, ChatBridge led its peers but achieved only 33.4% accuracy, far behind its human counterparts, highlighting the current neglect of cross-modal alignment by open-source audio-video MLLMs.
Performance of Proprietary MLLMs:
These models markedly outperform their open-source counterparts, achieving near-human proficiency on several tasks. Notably, Gemini achieved an average gap of just 15.2% to human results and performed well in four human-centric tasks: Human Emotion Perception, Person Recognition, Human Behavior Analysis, and Cross-Modal Speech-Visual Alignment.
Challenges with Time-Specific Tasks:
Many open-source video-MLLMs perform only marginally better than random chance, or even worse, on tasks requiring temporal-specific understanding. Tasks like Action at Specific Time, Time of Specific Action, Appearance Time Detection, and Audio-Visual Alignment Detection highlight deficiencies in accurately determining “video scene - event time” correlations, either bidirectionally or unidirectionally.
Cross-Modal Speech-Visual Alignment:
@@
Contemporary open-source models largely neglect intricate speech-to-lip movement alignment. Few have explicit architectural encoding linking audio and video modalities to facilitate such synergy, compounded by a scarcity of comprehensive datasets proposing coherent visual-audio lexical mapping. Consequently, these models struggle acutely in correlating speech to lip movements, though dedicated ASR models effectively address Speech Content Matching — a capability absent in tested open-source video-MLLMs.
@@ 4.3 Discussion on Speaker Emotion Recognition
A closer examination of these models reveals a tendency to incorrectly attribute speaker emotions as “surprise” or “shock”. This issue stems from frame sampling processes used by most models, where speaker frames depicting “mouth-opening” motions are misclassified. Figure 4 illustrates this phenomenon with the commonly used eight-frame fixed sampling [28, 53, 22, 51], resulting in erroneous emotional judgments due to temporal noisy information. Focused speaker-centric evaluations presented in Table 2 show noticeable accuracy depreciation, indicating the challenge of this task to these video MLLMs. On average, audio-visual models demonstrate slightly less decline compared to visual-only ones, as auditory data seemingly offers auxiliary emotional cues.
4.4 Class-Wise Analysis of Emotion Detection
For the Multi-modal Emotion Recognition in Conversation (MMERC) task, we visualized categorical emotion detection accuracy for zero-shot video emotion recognition settings, as shown in Figure 6.(a). Different MLLMs display variability in their effectiveness at recognizing diverse emotion types, with “neutral” emotions consistently yielding the highest accuracy. On average, these MLLMs show lower accuracy in identifying negative emotions such as “anger,” “disgust,” “sadness,” and “fear” compared to “neutral,” “joy,” and “surprise.” This suggests that MLLMs may be better at recognizing neutral or positive emotions or are more inclined to select neutral or positive emotions as their answers. Furthermore, as illustrated in Figure 6.(b), these MLLMs show varying capabilities in handling emotion polarities in the Human Emotion Recognition task, exhibiting higher precision for positive emotions, while neutral and negative emotional identification aligns closely in accuracy.
@@ 5 Conclusion
We present HumanVBench to address the pressing need for improved assessment of human-centric video understanding in MLLMs. By incorporating extensive evaluation dimensions through 17 fine-grained tasks, HumanVBench provides a systemic view into both successes and critical shortcomings of video MLLMs, particularly in emotion perception, human behavior analysis, and speech-visual alignment. Experimental findings across over ten leading video MLLMs illustrate that while proprietary models approach human accuracy in some tasks, substantial advancements are required, particularly in cross-modal domains where alignment between speech and visual elements proves challenging.
@@ 6 Overview
In the appendix, we first provide additional evaluation details in Section 7, followed by the detailed definition and examples for the 17 tasks of HumanVBench in Section 8. Then, we present implementation specifics of each operator for the proposed Human-Centric Annotation Pipeline, illustrated by an example of the annotation process in Section 9. Finally, we introduce the construction details of all tasks and the work of the human annotators in Section 10.
@@ 7 Model Evaluation Implementation
Prompt.
In order to facilitate the statistical model to answer the results, following common practices used in MLLM evaluations [18, 12], we adopt the following prompt to guide the MLLM to output option letters: “ Select the best answer to the following multiple-choice question based on the video. Respond with only the letter of the correct option. <Question-choices> Only answer best answer’s option letter. Your option is: ”.
Evaluation Environments.
All evaluation experiments for open-source models were conducted on a single NVIDIA L20 GPU with an inference batch size of 1.
Baseline Configurations and Runtime Statistics.
Table 3 shows the scale, parameter settings, and costs (including memory usage and end-to-end testing time) for each model on HumanVBench. The testing is repeated five times with different random seeds, and the results for all open-source models in Table 3 are based on the averages of these five runs. All hyperparameter settings follow the default configurations of these open-source works.
@@
Model Time (min) top_p top_k num_beams temp. VRAM
Chat-UniVi (7B) 44.7 1 50 1 0.2 14G
CogVLM2-Video (8B) 43.7 0.1 1 1 0.2 26G
Video-LLaVA (7B) 58.4 1 50 1 1 37G
LLaVA-OneVision (7B) 58.7 1 50 1 1 33G
PLLaVA (7B) 44.0 0.9 50 1 1 18G
ShareGPT4Video (8B) 55.2 0.9 50 1 1 17G
Otter-V (7B) 75.1 1 50 3 1 17G
VideoChat2-IT (7B) 62.9 0.9 50 1 1 18G
Video-LLaMA (7B) 48.4 1 50 2 1 22G
Video-LLaMA2.1 (7B) 49.3 0.9 50 1 0.2 20G
ImageBind-LLM (7B) 94.2 1 50 1 1 22G
ChatBridge (13B) 34.8 1 50 1 0.2 28G
OneLLM (7B) 146.2 0.75 50 1 0.1 16G
GPT-4o 232.0 1 - - 1 API
Gemini-1.5-Pro 313.9 1 40 - 0.9 API

### v2

@@ 3.3.2 Question and Preliminary Answer Synthesis
This stage synthesizes questions and initial answers by processing the video with bounding box track information of the target person. For tasks focusing on facial expressions (Human Emotion Recognition and Emotion Temporal Analysis), we reconstruct the video using cropped face regions; for others, we add a red bounding box to create “marked videos”. Then, we design prompts for Video-MLLM to generate task-specific captions centered on the target individual. We find that video-MLLMs don’t always focus correctly on the person highlighted by the bounding box. To address this, we enhance attention by adding a visual cue of the target person’s appearance from Label-4 (Figure 2).
@@
After acquiring task-specific captions, we analyze the caption attribute distributions (e.g., sentiment distribution for emotion recognition) and filter materials for question generation based on the desired distribution. Using these captions, we leverage SOTA LLMs to formulate questions and preliminary answers according to task definitions.
@@ 3.3.3 Answer Optimization and Distractors Generation
To obtain the most suitable answer, we iteratively optimize the answer using three Video-MLLMs while generating three distractors. Each Video-MLLM first generates a candidate answer from the video and question, then selects a better answer between this candidate and the current optimal answer (initialized as the preceding preliminary answer), updating the optimal answer. The discarded answers serve as raw material for distractor generation. This process repeats three times, producing one optimal answer and three discarded answers. An LLM then modifies the discarded answers with task-specific disturbances, ensuring distractors are distinct from the correct answer.
By using MLLM-discarded answers to synthesize options, the distractors are more challenging compared to those directly generated from the question and answer, thus more rigorously testing the model’s discriminative ability.
@@ 3.3.4 Manual Verification and Correction
Given the limitations of the annotation and question-answer generation models, errors in the multiple-choice questions may occur. To address this, we conduct manual validation as a final step. Specifically, we shuffle the options of each generated question and ask verifiers to select the best answer. If their choice matches the pipeline’s automatic selection, we confirm the question aligns with human cognition. If not, we re-examine and determine whether to revise the correct answer. If no suitable option is found, verifiers provide the correct answer, which becomes the definitive answer for the question. While this process involves some manual effort, it is significantly less labor-intensive and costly than generating questions and answers from scratch.
The Distractor-Included QA Synthesis Pipeline automates the tedious process of creating questions and options, shifting the human role from question creators to quality inspectors. Notably, we’ve decomposed the generation process into simpler sub-tasks, each manageable by open-source (M)LLMs. While these models are less accurate than their commercial counterparts, their inherent noise enhances the cognitive complexity of distractors. After human review and correction, this approach can still ensure high question quality while reducing labor costs. The pipeline is highly transferable, enabling its adaptation for designing a broader range of descriptive questions and choices.
3.4 Post-Processing
To address the prevalent issue of answer leakage [10] in multimodal evaluation datasets, we adopt the approach proposed in [10]. Specifically, we test all evaluated models without visual input and remove frequently correct QAs (approximately 6%) to ensure random accuracy. This strategy preserves the visual relevance of the questions while effectively mitigating the risk of answer leakage. Finally, HumanVBench includes 2116 problem instances. Instance counts for each task are detailed in Figure 1, and additional statistics are available in Appendix Section 7.
Figure 4: Two examples of 8-frame speaker videos sampled at equal intervals in the emotion recognition task, along with the responses from different MLLMs.
Models Chat-UniVi LLaVA-OneVision Intern VL2 Share-GPT4-Video Video Chat2-IT Qwen-VL2 LLaVA-Video Video-LLaMA Video-LLaMA-2.1 Chat Bridge GPT-4o Gemini-1.5-Pro Avg. Dec.
Emo. Acc. 29.2 36.5 35.4 36.5 25.0 41.7 40.6 25.0 32.3 27.1 54.1 57.3 36.7
Speaker Emo. Acc. 24.3$\downarrow$-4.9 34.3$\downarrow$-2.2 27.1$\downarrow$-8.3 32.8$\downarrow$-3.7 21.4$\downarrow$-3.6 37.1$\downarrow$-4.6 37.1$\downarrow$-3.5 22.9$\downarrow$-2.1 31.4$\downarrow$-0.9 21.4$\downarrow$-5.7 51.6$\downarrow$-2.5 54.0$\downarrow$-3.3 32.9$\downarrow$-3.8
Table 2: Accuracy of Video-MLLMs on Emotion Recognition (ER) for the full dataset and the subset where target individuals are in the speaking state. Current video-MLLMs struggle with emotion recognition for speaking individuals in HumanVBench.
@@ 4.1 Experimental Settings
We meticulously select 20 SOTA open-source video MLLMs. These included both visual-only models such as Chat-UniVi [25], CogVLM2-video [21], LLaVA-One-Vision [28], PLLaVA-7b [53], ShareGPT4Video [11], Otter-V [27], VILA [35], VideoChat [29], InternVL series [14, 15], Qwen-VL series [48, 4] and VideoLLaMA3 [57] and audio-visual models capable of analyzing both visual and audio inputs, such as the Video-LLaMA series [58, 16] and generalist MLLMs like ImageBind-LLM [19], ChatBridge [61], and OneLLM [20]. We also evaluated commercial models GPT-4o [41] and Gemini-1.5-Pro [47]. All QAs were framed as multiple-choice questions (N choose 1, with N varying across different test samples), reporting both accuracy and the performance of random guesses and graduate-level humans for reference. More implementation details can be found in the Appendix Section 9.
@@ 4.2 Main Results
Table 1 summarizes the performance evaluation results for our benchmark, revealing several key insights from analyzing vision-only tasks (emotion perception, person recognition, and behavior analysis tasks) and cross-modal tasks (speech-visual alignment tasks).
4.2.1 Performance in Vision-Only Tasks
Open-Source Video-MLLMs: Generally, these models still show a noticeable gap compared to human-level performance, particularly in Emotion Perception tasks, highlighting the need for further development to bridge this gap. However, some models have demonstrated superior performance in certain tasks, rivaling or even surpassing commercial models, such as LLaVA-OneVision in AR and VideoLLaMA3 in T2H, ATD, and AST. Among open-source models, VideoLLaMA3 leads in video human understanding, and achieves a modest 58.4% mean accuracy across 12 vision-only tasks. Although it still lags behind human performance (88.2%), it has surpassed the commercial model GPT-4o (51.7%) and is close to Gemini-1.5-pro (65.5%), demonstrating the immense potential of open-source models.
Proprietary MLLMs: Overall, these models outperform their open-source counterparts in most tasks, achieving near-human proficiency in several tasks. Notably, Gemini delivers superior performance compared to other models, while GPT-4o struggles in tasks like AR, T2H, and ATD. As a result, GPT-4o’s average performance in Emotion Perception and Person Recognition lags behind several leading open-source models.
4.2.2 Performance in Speech-Visual Alignment Tasks
Lip-Reading Ability of Visual-Only MLLMs For AVSM and ASD tasks, many videos include dubbing or multi-speaker conversations with single-speaker audio, making lip movements alone insufficient for accurate responses. Therefore, when answering these two tasks, visual-only models essentially degrade to speech action recognition but can still get some questions correct. As shown in Table 1, these MLLMs show some lip-reading ability, with QwenVL2 and InternVL2 outperforming other models. However, for the SCM task, almost all models perform at a random level, indicating that current models lack precise lip-reading (lip translation) ability. Future models could focus on enhancing lip-reading capabilities.
Error Type Error Percentage
Misidentifying emotion as Surprise 41%
Overinterpretation of neutral emotions 29%
Emotion Polarity Mistakes 29%
Table 3: Error Analysis of Emotion Recognition in Gemini
Video-MLLMs ATD AST TSA
Chat-UniVi 17.5 ($\uparrow$2.3) 40 ($\uparrow$26) 14.1 ($\uparrow$7.4)
CogVLM2 40.8 ($\uparrow$0.8) 35 ($\downarrow$1) 10.7 ($\downarrow$3.4)
VILA 38.3 ($\uparrow$18.1) 33 (–) 48 ($\downarrow$3.4)
Video-LLaVA 36.7 ($\uparrow$5) 25 ($\downarrow$2) 45.8 ($\uparrow$5.1)
LLaVAOneVision 10 (–) 41 ($\uparrow$4) 47.5 (–)
InternVL2 31.7 ($\uparrow$10.9) 45 ($\uparrow$7) 52.5 ($\uparrow$18.7)
Video-LLaMA 24.2 ($\uparrow$4.2) 25 ($\uparrow$7) 9 (–)
Video-LLaMA2.1 24.2 ($\uparrow$7.5) 25 ($\downarrow$2) 29.4 ($\downarrow$3.9)
ChatBridge 17.5 ($\uparrow$8.8) 40 ($\downarrow$1) 14.1 ($\downarrow$2.8)
Average 26.8 ($\uparrow$5.9) 34.3 ($\uparrow$4.4) 30.1 ($\uparrow$2.0)
Table 4: Timestamp integration effect on Video-MLLMs in Appearance Time Detection (ATD), Action at Specific Time (AST) and Time of Specific Action (TSA).
@@ 4.2.2 Performance in Speech-Visual Alignment Tasks
Audio-Video MLLMs: Despite handling both audio and visual inputs, these open-source MLLMs perform near-randomly across Speech-Visual Alignment tasks, particularly in AVSM and ASD, where they lag behind visual-only MLLMs. On one hand, this may stem from deficiencies in the models’ ability to visually interpret lip movements (see Appendix Section 8 for extended experiments); on the other, contemporary open-source models largely neglect intricate speech-to-lip movement alignment. Few have explicit architectural encoding linking audio and video modalities to facilitate such synergy, compounded by a scarcity of comprehensive datasets proposing coherent visual-audio lexical mapping. Consequently, these models struggle acutely in correlating speech to lip movements, though dedicated ASR models effectively address Speech Content Matching — a capability absent in tested open-source video-MLLMs.
@@ 4.3 Discussion on Speaker Emotion Recognition
A detailed analysis reveals that models frequently misattribute speaker emotions as “surprise” or “shock.” For example, Gemini-pro-1.5, the top performer in Emotion Recognition, exhibits this tendency, with surprise misidentification accounting for 40% of its errors (Table 2). This issue primarily stems from frame sampling processes, where frames depicting “mouth-opening” motions are misclassified. Figure 4 illustrates this phenomenon using the commonly adopted eight-frame fixed sampling approach [34, 60, 28, 58], which introduces temporal noise and leads to erroneous emotional judgments. The speaker-centric evaluations presented in Table 2 further confirm a consistent decline in accuracy across all models, underscoring the inherent challenges of this task for video MLLMs.
4.4 Impact of Timestamp Inclusion on Time-Specific Tasks
Table 1 shows that many open-source video-MLLMs struggle with time-specific tasks—like Appearance Time Detection, Action at Specific Time, and Time of Specific Action—due to challenges in establishing “video scene-event time” correlations. Analysis shows that only a few models, including top performers like VideoLLaMA3 and LLaVA-Video, explicitly incorporate textual timestamps as part of their built-in prompt processing from sampled video frames. To explore the impact of timestamp integration, we added timestamps (e.g., “You have read N frames, corresponding to seconds in the video as: […]”) to models without native support. As shown in Table 4, this intervention generally improved temporal reasoning accuracy, though its effectiveness varied across models, likely due to limited exposure to timestamp-related data, reflecting the models’ inherent capabilities and potential.
@@ 5 Conclusion
We present HumanVBench to address the pressing need for improved assessment of human-centric video understanding in MLLMs. By incorporating extensive evaluation dimensions through 16 fine-grained tasks, HumanVBench provides a systemic view into both successes and critical shortcomings of video MLLMs, particularly in emotion perception and speech-visual alignment. Experimental findings across over twenty leading video MLLMs illustrate that while proprietary models approach human accuracy in some tasks, substantial advancements are required, particularly in cross-modal domains where alignment between speech and visual elements proves challenging.
@@ 6 Overview
In the appendix, we first We provide more benchmark statistics in Section 7, then the modality ablation experiments in VideoLLaMA2 in Section 8, followed by additional evaluation details in Section 9, followed by the detailed definition and examples for the 16 tasks of HumanVBench in Section 10. Then, we present implementation specifics of each operator for the proposed Human-Centric Annotation Pipeline, illustrated by an example of the annotation process in Section 11. Finally, we introduce the construction details of all tasks and the work of the human annotators in Section 12.
7 More Statistics of HumanVBench
HumanVBench focuses on short video understanding, specifically videos with a duration of 10 seconds or less. It includes a total of 2116 question instances, with the specific number for each task indicated in 1. The total video duration amounts to 4.7 hours and demonstrates a variety of people, scenes, and video shooting styles, as shown in 6.
Figure 6: The distribution of the number of people, scenes, and video shooting styles in HumanVBench
8 Modality Ablation in VideoLLaMA2
Models Input Modal Human Emotion Perception Person Recognition Human Behavior Analysis Speech-Visual Alignment
ER ETA AR EIC Avg T2H H2T HC ATD Avg BTA BCA AST TSA Avg AVSM ASD AVAD SCM Avg
Random 25 25.0 25.0 25.0 22.9 27.9 25.0 23.1 25.0 25.3 25.0 25.0 25.0 20.0 23.8 42.8 23.6 33.3 25.0 31.2
Video-LLaMA2.1-AV A, V 32.3 28.7 47.0 21.7 32.4 34.6 47.9 41.5 16.7 35.2 47.5 43.3 27.0 33.3 37.8 44.0 31.6 32.1 23.7 32.9
Video-LLaMA2.1-AV V 27.1 31.6 45.1 17.2 30.3 36.9 46.8 36.8 15.0 33.9 43.2 50.7 31.0 29.9 38.7 43.4 29 32.8 18.8 31.0
VideoLLaMA2-7B-16F V 32.3 30.9 34.8 23.7 30.4 37.4 51.1 36.8 15.8 35.3 53.4 50.7 32.0 49.2 46.3 47.4 38.1 32.1 15.4 33.3
Table 5: The performance of VideoLLaMA2’s vision-only version (VideoLLaMA2-7B-16F), and the audio-visual version (VideoLLaMA2.1-AV), on HumanVBench, based on different modal inputs (A for Audio, V for Visual).
Despite audio-visual MLLMs processing audio data, they perform at random-guess levels on AVSM and ASD tasks, underperforming relative to many vision-only models that rely solely on lip movement analysis. This raises the question: does the poor performance stem from limitations in visual analysis (e.g., lacking lip-reading ability) or from the interference of audio input? To explore this, we conducted ablation experiments using the VideoLLaMA2 model series, chosen for its open-source availability of both vision-only and audio-visual variants.
As shown in the Table 5, VideoLLaMA2-7B-16F (vision-only) exhibits only a slight advantage over Video-LLaMA2.1-AV (audio-visual) on AVSM and ASD tasks, yet still lags far behind vision-only models such as VideoLLaMA3 and InternVL2.5 (Table 1). This indicates that VideoLLaMA2-7B has inherently poor lip-reading capability, which further implies that the audio-visual variant (Video-LLaMA2.1-AV) also suffers from limited visual lip-reading ability. Such limitations constrain its upper-bound performance in speech-visual alignment tasks. On the other hand, Video-LLaMA2.1-AV shows no significant performance advantage when utilizing audio information compared to its vision-only counterpart. This suggests that vocal information is not effectively leveraged, likely due to insufficient speech parsing capability in video MLLMs and inadequate understanding of cross-modal associations between audio and visual content.
@@ 9 Model Evaluation Implementation
Prompt. In order to facilitate the statistical model to answer the results, following common practices used in MLLM evaluations [24, 18], we adopt the following prompt to guide the MLLM to output option letters: “ Select the best answer to the following multiple-choice question based on the video. Respond with only the letter of the correct option. <Question-choices> Only answer best answer’s option letter. Your option is: ”. Evaluation Environments. All evaluation experiments for open-source models were conducted on a single NVIDIA L20 GPU with an inference batch size of 1.
Baseline Configurations and Runtime Statistics. Table 6 shows the scale, parameter settings, and costs (including memory usage and end-to-end testing time) for each model on HumanVBench. All hyperparameter settings follow the default configurations of these open-source works.
Model Time (min) top_p top_k num_beams temp. VRAM
Chat-UniVi (7B) 35 1 50 1 0.2 14G
CogVLM2-Video (8B) 34 0.1 1 1 0.2 26G
Video-LLaVA (7B) 46 1 50 1 1 37G
LLaVA-OneVision (7B) 46 1 50 1 1 33G
PLLaVA (7B) 35 0.9 50 1 1 18G
ShareGPT4Video (8B) 44 0.9 50 1 1 17G
Otter-V (7B) 60 1 50 3 1 17G
VideoChat2-IT (7B) 50 0.9 50 1 1 18G
InternVL2 (7B) 44 1 50 1 1.0 21G
InternVL2.5 (7B) 31 1 50 1 1.0 21G
Qwen2-VL (7B) 43 0.001 1 1 0.1 22G
Qwen2.5-VL (7B) 33 0.001 1 1 0.1 22G
LLaVA-Video (7B) 173 0.8 20 1 0.7 36G
Video-LLaMA3 (7B) 86 0.8 20 1 0.7 21G
Video-LLaMA (7B) 38 1 50 2 1 22G
Video-LLaMA2.1 (7B) 27 0.9 50 1 0.2 23G
ImageBind-LLM (7B) 75 1 50 1 1 22G
ChatBridge (13B) 27 1 50 1 0.2 28G
OneLLM (7B) 116 0.75 50 1 0.1 16G
GPT-4o 185 1 - - 1 API
Gemini-1.5-Pro 250 1 40 - 0.9 API